### v1

# Parallel Cluster-BFS and Applications to Shortest Paths

Abstract
Breadth-first Search (BFS) is one of the most important graph processing subroutines, especially to compute the unweighted distance. Many applications may require to run BFS from multiple sources. Sequentially, when running BFS on a cluster of nearby vertices, a known optimization is to use bit-parallelism. Given a subset of vertices with size $k$ and the distance between any pair of them is no more than $d$, BFS can be applied to all of them in a total work of $O(dk/w+1)$, where $w$ is the length of a word in bits. We will refer to this approach as cluster-BFS (C-BFS). Such an approach has been studied and shown effective both in theory and in practice in the sequential setting. However, it remains unknown how this can be combined with thread-level parallelism for C-BFS.
In this paper, we focus on designing efficient parallel C-BFS based on BFS to answer unweighted distance queries. Our solution combines the strengths of bit-level parallelism and thread-level parallelism, and achieves significant speedup over the plain sequential solution. We also apply our algorithm to real-world applications. In particular, we identified another application (landmark-labeling for the approximate distance oracle) that can take advantage of parallel C-BFS. Under the same memory budget, our new solution improves accuracy and/or time on all the 18 tested graphs.
We released our code [53] and graph instances used in the experiments [54].
1   Introduction
Breadth-First Search (BFS) is one of the most important graph processing subroutines. Given a graph $G=(V,E)$ and a vertex $s\in V$, BFS visits all vertices in $V$ in increasing order of (hop) distance to $s$. BFS can be used for many purposes. One of the most common use scenarios for BFS is to compute the unweighted distance from the source. In this paper, we focus on designing efficient parallel approaches based on BFS to answer unweighted distance queries. Throughout the paper, we use $n=|V|$ and $m=|E|$, and use “distance” to refer to the hop distance on an unweighted graph.
Many applications may require to run BFS from multiple sources. Examples of this are using BFS for low-diameter decomposition [37], all-pairs shortest paths (APSP), or oracles for exact or approximate APSP. A key observation is that bit-parallelism [20] can be used effectively when running BFS on a cluster of nearby vertices [20, 3]. Given a subset of vertices with size $k$ and the distance between any pair of them is no more than $d$, BFS can be applied to all of them in total $O(dk/w+1)$ work (number of operations), where $w$ is the length of a word in bits [20]. Since a machine word must hold at least $\Omega(\log n)$ bits to store a pointer, this means $w=\Omega(\log n)$. We will refer to this approach as cluster-BFS (C-BFS), and present more details in Sec. 3. In 2012, Chan [20] first observed this and used this idea to develop an all-pair shortest-path algorithm that runs in $O(mn/w)$ work (when $m=\Omega(n\log n\log\log\log n))$. In addition to saving time, C-BFS also saves space: it only uses $O(d)$ words per $w$ vertices instead of a word (or at least enough bits to store a distance) per vertex as in standard BFS. In 2013, Akiba et al. [3] used this idea in the exact two-hop distance oracle but only considered the special case for $d=2$ (a star-shaped cluster: a vertex and its neighbors). We refer to this algorithm as the AIY algorithm. Both of the previous papers focus on the sequential setting.
While the C-BFS with bit-level parallelism has shown effective in sequential settings, surprisingly, we know of no previous work studying the (thread-level) parallelism on it. BFS is one of the most well-studied parallel graph processing problems, and state-of-the-art solutions have been highly optimized using techniques such as direction optimizations [8, 45]. To be practical, any C-BFS would have to compete with these. Our goal is to develop efficient C-BFS with high parallelism such that it 1) achieves the same level of parallelism as the standard parallel BFS, with additional benefits by using clusters, 2) supports a clean interface that is flexible for different parameter settings (i.e., varying $d$ and $k$), and 3) facilitates various real-world applications. In this paper, we provide a systematic study of parallel C-BFS and achieve all three goals above.
To achieve high performance, we design an efficient parallel algorithm. Our algorithm is work-efficient (i.e., it has the same asymptotical work as the sequential counterpart). It has the same span as regular parallel BFS algorithms (e.g., [45]), which $\tilde{O}(D)$ for graph diameter $D$, and thus is best suited for small-diameter graphs, such as social networks, computer networks, or web graphs. Our algorithm uses the directional optimization that has been shown to be useful for parallel BFS. By doing this, our algorithm achieves the strengths of both bit-level and thread-level parallelism: it has high parallelism as in the state-of-the-art parallel BFS algorithms and obtains additional performance gain by using bit-level parallelism.
To achieve a flexible interface, we designed our algorithm to work for general $k$ and $d$ and can be easily integrated into different applications using user-defined parameters.
To understand how C-BFS can facilitate real-world applications, we study two Distance Oracle (DO) techniques that can benefit from C-BFS: the exact DO as in [3] and the Landmark Labeling (LL) for an approximate DO. As far as we know, our work is the first to use C-BFS to accelerate LL.
We implemented our C-BFS algorithm and the two applications. We compared our algorithm with multiple baselines to study the performance gain in depth. We tested 18 graphs with various types and sizes on a 96-core machine. Compared to standard sequential BFS, our algorithm employs both bit-parallelism (on clusters) and thread-level parallelism (along with optimizations used in parallel BFS) to improve performance. In the simplest setting where $k=64$ and $d=2$, the combination of them enables up to 1119$\times$ speedup (500$\times$ on average) compared to the plain sequential BFS, where bit-level and thread-level parallelism each contributes about 20$\times$ speedup. Interestingly, by comparing with the performance of existing work (Ligra [45], where only thread-parallelism is used, and AIY [4], where only bit-level parallelism is used), we observed that both bit-level parallelism and thread-level parallelism and work very well in synergy. Each of them still fully contributes to the performance when the other presents, achieving the same level of improvement as when used independently (see Fig. 1). Therefore, we believe our work on an efficient implementation combining thread-level and bit-level parallelism fills the gap in the existing study of both C-BFS and parallel BFS.
We also studied the performance of our C-BFS with different parameters. Typically, $k$ is set to be $\Theta(w)$, and the work (and space) is proportional to $d$. Our result shows that the running time increases almost linearly with value of $d$, especially when $d$ is small. This explains why existing work (e.g., [4]) tends to choose the smallest $d=2$ case in real-world applications.
Applying our algorithm also gives significant improvement to the aforementioned applications. For the 2-hop distance oracle, our parallel implementation outperforms the sequential AIY algorithm [3] by 9–36$\times$, and can process much larger graphs than the AIY algorithm. For landmark labeling (LL), with a fixed memory budget, C-BFS improved regular LL in either accuracy or preprocessing time on all 18 tested graphs, and improved both on 14/18 graphs. This is due to the saving in space allowing more landmarks to be used for C-BFS. We observed that using $d=2$ achieved better overall performance in accuracy, time and space. Due to the page limit, we present more results in the Appendix.
Figure 1: Performance comparison with existing work. We test the running time of BFSs from a cluster of 64 vertices. The baselines are Ligra [45] that only uses thread-level parallelism and AIY’12 [4] that only uses bit-level parallelism. The numbers are geometric means across 18 graphs. Full results are shown in Tab. 2 and Fig. 3.
2   Preliminaries
Notations.   Let $G=(V,E)$ be an unweighted graph. We use $n=|V|$, $m=|E|$, and use $D$ to denote the diameter of the graph. Let $N(v)=\{u\in V~{}|~{}(v,u)\in E\}$ be the set of neighbors of vertex $v\in V$. In directed graphs, $N^{+}(v)$ and $N^{-}(v)$ represent outgoing and incoming neighbors, respectively. We use $\delta(u,v)$ to denote the shortest distance between $u$ and $v$. We assume machine word size $w=\Omega(\log n)$, such that the vertex and edge IDs are within constant words. Let $S=\{s_{1},s_{2},...,s_{k}\}$ represent a cluster, where $k$ is the cluster size. Let $d$ be the diameter (maximum distance between any pair) of the cluster. We summarize the notations in Tab. 1.
Computational Model.   We use the binary fork-join parallelism [23, 10], with work-span analysis [14, 28]. We assume a set of threads that access a shared memory. A process can fork two child software threads to work in parallel. When both children complete, the parent process continues. A parallel for-loop can be simulated by recursive forks in logarithmic levels. The work of an algorithm is the total number of instructions, and the span is the length of the longest sequence of dependent instructions. We can execute the computation using a randomized work-stealing scheduler [14, 27] in practice.
[…]
Parallel BFS.   We briefly review parallel BFS, because it is one of our baselines and some of the concepts are also used in our cluster-BFS. Parallel BFS starts from a single source $s\in V$ (high-level idea in Alg. 3). The algorithm maintains a frontier of vertices to explore in each round, starting from the source, and finish in $D$ rounds. In round $i$, the algorithm processes (visits their neighbors) of the current frontier $\mathcal{F}_{i}$, and puts all their (unvisited) neighbors to the next frontier $\mathcal{F}_{i+1}$. If multiple vertices in $\mathcal{F}_{i}$ attempt to add the same vertex to $\mathcal{F}_{i+1}$, a compare_and_swap is used to guarantee that only one will succeed.
One widely-used optimization for parallel BFS is the directional optimization [8, 45]. At a high level, when the frontier size $|\mathcal{F}_{i}|$ is large, the algorithm will not process $\mathcal{F}_{i}$, but instead visit each unprocessed vertex $v$, and determine if $v$ has an incoming neighbor in $\mathcal{F}_{i}$. If so, $v$ will be put in $\mathcal{F}_{i+1}$. Such an optimization is observed to be effective, especially on small-diameter graphs. We present more details in Sec. A.
$G=(V,E)$
: the input graph. $n=|V|$ and $m=|E|$.
$S=\{s_{1},...,s_{k}\}$
: the source cluster for the BFS.
$k$
: the cluster size, i.e., $k=|S|$.
$d$
: the diameter of the cluster.
$w$
: the length of a word in bits. $w=\Omega(n)$.
$D$
: the diameter of the graph.
$\delta(u,v)$
: the shortest distance between $u$ and $v$.
Table 1: Notations in the paper.
3   Parallel Cluster-BFS
Cluster-BFS (C-BFS) runs BFS from a cluster of sources $S\subseteq V$. If the sources have diameter $d$ (maximum distance between any pair), then all distances from $S$ to any vertex $v\in V$ will differ by at most $d$. This means that if we run a set of BFSs from $S$, synchronously, every $v\in V$ will appear in at most $d+1$ consecutive frontiers. Cluster-BFS takes advantage of this by representing all the sources in $S$ that can visit a given vertex, on a given round, as a vector of booleans (bits). In this way, each vertex will be visited at most $d+1$ times instead of $|S|$ times if all searches are performed separately. We will explain more details in section 3.1. Importantly, if the bit-vector fits in $O(1)$ words, the distances of all $|S|$ sources can be handled (propagated from a vertex in the current frontier to a neighbor) with $O(1)$ bitwise logical operations. Since a machine word must hold at least $\Omega(\log n)$ bits (so it can represent a pointer), this means the algorithm can save a factor of $\Omega(\log n/d)$ work.
It appears that Chan first described this idea [20], but he did not go into any details of the implementation, but just saying that this is possible. Akiba et al. [3] later showed a concrete implementation based on this idea, with the limitations that it is sequential and only works on the cluster of a star with $d=2$ (a center vertex and its neighbors). To the best of our knowledge, there has been no previous work on developing a parallel implementation of the cluster-BFS algorithm. Indeed this is pretty challenging to achieve high performance given how BFS has been widely studied with numerous optimizations both sequentially and in parallel. In this paper, we propose our algorithm, given as Alg. 1, which is an efficient parallel cluster-BFS implementation with a general interface and low coding effort. In the following sections, we will first introduce the bitwise representation to maintain the distances in cluster-BFS and then give our parallel cluster-BFS algorithm.
3.1   Cluster Distance Representations
Given a set $S$ of source vertices with diameter $d$, cluster-BFS computes a compact representation of the shortest distance from every source in $S$ to every vertex in $V$. The idea of cluster-BFS is based on the following fact.
Fact 3.1
On an unweighted graph, if the distance betweeen vertex $s_{1}$ and $s_{2}$ is $d$, then for any vertex $v\in V$, $|\delta(s_{1},v)-\delta(s_{2},v)|\leq d$.
For example, if $s_{1}$ and $s_{2}$ are neighbors, the distances from a vertex $v$ to them can differ by at most 1. We can further extend Fact 3.1 to Cor. 3.1, which says if a cluster of vertices $S$ has diameter $d$, the distances from $v$ to vertices in $S$ differ by at most $d$.
Corollary 3.1
On an unweighted graph, given a set $S$ of vertices with diameter no more than $d$, for any vertex $v\in V$, we have
$\max_{s\in S}\delta(s,v)-\min_{s\in S}\delta(s,v)\leq d$
Therefore, for each vertex $v$, we can classify the sources in $S$ by their distances to $v$. Let $\Delta_{v}=\min_{s\in S}\delta(s,v)$ be the smallest distance from any source in $S$ to $v$. According to Cor. 3.1, the distance between $v$ and any $s\in S$ must be in range $[\Delta_{v},\Delta_{v}+d]$. This divides all vertices in $S$ in $d+1$ different subsets based on their distances to $v$. Let ${S}_{v}[i]$ be the subset of sources in $S$ that has a distance to $v$ as $\Delta_{v}+i$. More formally,
${S}_{v}[i]=\{s\in S~{}|~{}\delta(s,v)=\Delta_{v}+i\}$
Then for a vertex $v$, the distances between $v$ and all sources in $S$ can be represented by the $(d+2)$-tuple $\langle{S}_{v}[0..d],\Delta_{v}\rangle$, which we call the cluster distance vector of $v$ to $S$.
Note that if $|S|=w$, we can use one word of bit-vector to represent any subset of $S^{\prime}\subseteq S$: bit $i$ is 1 iff. the $i$-th element in $S$ is also in $S^{\prime}$. We call such a bit-flag representation of a subset of $S$ a bit-subset. In this way, a cluster distance vector only takes $d+1$ words for bit-subsets and one byte for distance to store the shortest distances from $v$ to $|S|$ sources.
Figure 2: Illustration of bitwise representation. The batch set $S$ is $\{A,B,C,D\}$. 4-bit bit-subsets are used to represent subsets of $S$. $\Delta_{v}$ is the smallest shortest distance from any vertex in $S$ to $v$. The subset ${S}_{v}[i]$ is defined as $\{s\in S|\delta(s,v)=\Delta_{v}+i\}$.
An illustration for the bit-subset and cluster distance vector is shown in Fig. 2. In this example, $S$ is the set $\{A,B,C,D\}$, and the diameter of the subgraph is $d=2$. We need subsets ${S}_{v}[0..2]$ for each vertex $v$, which are represented by the bit-subsets, each with four bits. $A$ to $D$ are represented by the four bits from left to right. From the cluster distance vector, we can recover the shortest distance from all the sources $s\in S$ to each vertex $v\in V$ by the fact that each source in ${S}_{v}[i]$ has distance $\Delta_{v}+i$ to $v$. For example, for vertex $F$, ${S}_{F}[1]$ is 1010, which represents the subset $\{A,C\}$, we can infer $\delta(B,A)=\delta(B,C)=\Delta_{F}+1=2$.
The main idea of cluster-BFS is to use bitwise operations on these bit-subsets to quickly compute the union/intersection of the sets, so that we can use the cluster distance vector of $v$ to compute the cluster distance vectors of its neighbor $u$ in constant time. In the following, we elaborate on our parallel cluster-BFS algorithm.
Input:
A graph $G=(V,E)$, a cluster $S\subseteq V$ with diameter $d$
Output:
cluster distance vectors $\langle{S}_{v}[0..d],\Delta_{v}\rangle$ for all $v\in V$.
Maintains :
$i$: the current round number, initialized to 0
$S_{seen}[\cdot],S_{next}[\cdot]$: array of bit-subset for each $v\in V$
$r[v]$: the lastest round $v$ is in the frontier
$\mathcal{F}_{i}$: frontier vertices in round $i$
// Initialization
1 ParallelForEach $v\in V$ do
2       $S_{seen}[v]\leftarrow\emptyset$, $S_{next}[v]\leftarrow\emptyset$
3       $\Delta_{v}\leftarrow\infty$
4       $r[v]\leftarrow\infty$
5for $s\in S$ do  $S_{seen}[s]\leftarrow\{s\}$
6 $i\leftarrow 0$
7 $\mathcal{F}_{0}\leftarrow S$
// Traversing
8 while $\mathcal{F}_{i}\neq\emptyset$ do
9       ParallelForEach $u\in\mathcal{F}_{i}$ do
10             $S_{new}\leftarrow S_{next}[u]\setminus S_{seen}[u]$
11             if $\Delta_{u}=\infty$ then  $\Delta_{u}\leftarrow i$
12             ${S}_{u}[i-\Delta_{u}]\leftarrow S_{new}$
13             $S_{seen}[u]\leftarrow S_{seen}[u]\cup S_{new}$
14      ParallelForEach $u\in\mathcal{F}_{i}$  do
15             ParallelForEach $v\in N(u)$ and $i-\Delta_{v}<d$  do
16                   if Fetch_And_Or($S_{next}[v],S_{seen}[u]$)
17                         if compare_and_swap$(r[v],r[v],i)$
18                               $\mathcal{F}_{i+1}\leftarrow\mathcal{F}_{i+1}\cup\{v\}$
19      $i\leftarrow i+1$
20
return $\langle{S}_{v}[1..d],\Delta_{v}\rangle$ for all $v\in V$
Algorithm 1 Cluster-BFS search from $S$
3.2   Our Parallel Algorithm
In this section, we introduce our parallel algorithm to compute the cluster distance vector for all vertices given a source cluster $S$, which is $\langle{S}_{v}[0..d],\Delta_{v}\rangle$ for all $v\in V$. The pseudocode of our cluster-BFS is shown in Alg. 1.
Our cluster-BFS algorithm is based on the following fact: if $u$ and $v$ are neighbors and there is a path from a source $s\in S$ to $u$ with length $i-1$, then there must exist a path from $s$ to $v$ with length $i$. In round $i$, $u$ records all the sources in $S$ that reach it in round $i$ by a bit-subset. When $u$ visits its neighbor $v$, $u$ propagates this bit-subset to $v$ by taking a bitwise Or operation with the bit-subset representing the vertices reaching $v$ in round $i+1$ (Alg. 1: Alg. 1). According to Cor. 3.1, all the vertices in $S$ will visit $v$ at least once during round $\Delta_{v}$ to round $\Delta_{v}+d$; in other words, all the vertices will be put into the frontier for $d+1$ times. Therefore, we need to record the number of times $v$ has been put in the frontier. When $v$ has been put in the frontier for $d+1$ times, since all sources in $S$ must have already visited $v$, we do not need to process $v$ anymore. Otherwise, we will process $v$ and put it to the next frontier since other sources in $S$ may visit $v$ in the future.
Our cluster-BFS algorithm is given in Alg. 1. In our algorithm, we use a boolean array $S_{seen}[\cdot]$ to store whether a vertex has been visited in all previous rounds by any sources in $S$, and $S_{next}[v]$ includes the vertex if it is also in the current frontier (i.e., visited by any vertex in the current round). We denote $\Delta_{v}$ as the first round that any vertex from $S$ touches vertex $v$. Alg. 1 has two stages: initialization and traversing.
Initialization   This step is relatively simple. We initialize the arrays of $S_{seen}[\cdot]$, $S_{next}[v]$, and $\Delta_{v}$. We use another, array $r[\cdot]$, to avoid duplication of vertices in the frontier. Later in the traversing stage, when multiple vertices want to add $v$ to the next frontier at the same time, only one can successfully set $r[v]$ to the current round number by atomic operation compare_and_swap (Alg. 1), and the successful vertex will put $v$ to the next frontier.
Traversing   At the beginning, the algorithm puts all the sources $s\in S$ into the first frontier $\mathcal{F}_{0}$. Then, we visit all vertices by frontiers. In each round, we process frontiers in two stages, where the first stage processes vertices and the second stage processes edges. In the first stage (Alg. 1 to Alg. 1), we first compute the sources that newly visited $u$ by $S_{new}\leftarrow S_{next}[u]\setminus S_{seen}[u]$ (Alg. 1). Note that the bit-subset $S_{new}$ contains sources whose distances are $i$, which is also ${S}_{u}[i-\Delta_{u}]$ (Alg. 1). Then we update $S_{seen}[u]$ to include newly visited vertices (Alg. 1), and set $\Delta_{u}$ to the current round number if it has not been set yet (Alg. 1). In the second stage (Alg. 1 to Alg. 1), we process the neighbor vertices of the current frontier that have not been visited for $d$ times already (Alg. 1). For an edge from $u\in\mathcal{F}_{i}$ to its neighbor $v$, we propagate the sources seen so far by $u$, $S_{seen}[u]$, to $v$ (Alg. 1). In general, if any source $s\in S$ visited $u$ in the previous round, $s$ should also visit $v$ in this round, and should be included in the $S_{next}[v]$ for $v$ in this round. If the $S_{next}[v]$ is changed, which means there are new sources visiting $v$, $v$ should be added to the next frontier. To avoid duplication in the next frontier, only the one that can successfully set $r[v]$ to $i$ (Alg. 1) by compare_and_swap will put $v$ to the next frontier. Note that we can further benefit from the directional optimization that is commonly used in parallel BFS. Additional details about the directional optimization are available in Sec. A.
The efficiency of the algorithm relies on using bit-operations to compute the union and difference of two bit-subsets, stated below.
Lemma 3.1
Given the bit-subset subsets $S_{1},S_{2}$ of a set $S$ with size $k$, we can compute the bit-subset representation of $S_{1}\cup S_{2}$, $S_{1}-S_{2}$ in $O(k/w+1)$ work and $O(\log(k/w)+1)$ span, where $w$ is the word length.
  • Proof.
    A bit-subset with $k$ bits needs $\lceil k/w\rceil$ words. These $\lceil k/w\rceil$ words can be processed in parallel. With the constant cost for bitwise or/not operations for a single word, the work and span for $S_{1}\cup S_{2}$, $S_{1}-S_{2}$ are $O(k/w+1)$ and $O(\log(k/w)+1)$.
We now show the cost analysis of the cluster-BFS algorithm.
Theorem 3.1
Given a set $S$ of $k$ vertices with diameter $d$, we can compute the cluster distance vector from $S$ to every vertex in $V$ in $O(dm(k/w+1))$ work and $O((D+d)\log n)$ span.
  • Proof.
    We will analyze the work and span for the traversal stage, which dominates the cost of the initialization stage. The traversal consists of two phases: the first phase processes vertices in the frontier, and the second phase processes edges from the frontier. The cost of processing a single edge and a vertex is asymptotically the same, as it involves applying a constant number of set operations, which results in $O(k/w+1)$ work and $O(\log(k/w)+1)$ span, as described in Lem. 3.1. Assuming $n<m$, the cost of traversal is primarily determined by edge processing. Therefore, the rest of the proof will focus on analyzing the cost of edge processing.
    For the work, since each vertex is in the frontier for at most $d$ times, each edge is processed at most $d$ times, leading to a total work of $O(dm(k/w+1))$. For the span, recall that $D$ is the diameter of the graph, and there are at most $D+d$ rounds in the traversal stage. The span for each round is $O(\log(k/w)+1+\log n)$, which accounts for the cost of generating $O(m)$ parallel tasks (costing $O(\log n)$ in the binary fork-join model) and the cost of processing a single edge (costing $O(\log(k/w)+1)$ as shown in Lem. 3.1). Since $k/w$ is smaller than $n$, the span for each round simplifies to $O(\log n)$. Thus, the total span for the $D+d$ rounds of the traversal stage is $O((D+d)\log n)$. Therefore, the total work and span for our parallel cluster-BFS are $O(dm(k/w+1))$ and $O((D+d)\log n)$, respectively.
If we take $k=O(w)$, such that each bit-subset fits within a constant number of words, the work simplifies to $O(dm)$ and the span remains $O((D+d)\log n)$, which matches the work and span of a single BFS. Since $w=O(\log n)$, this means that we can compute $O(\log n)$ more BFSs with asymptotically the same cost.
4   Applications
In this section, we show two applications that can benefit from our new parallel Cluster-BFS algorithm. Both applications are distance oracles (DO) for unweighted graphs. A distance oracle (DO) is an index designed to answer the shortest distances between two vertices on a graph. Although such a query can always be answered by computing the distance on the fly (e.g., running a BFS from one of the query vertices), it can be inefficient for applications requiring low latency or requesting multiple queries. A distance oracle aims to store certain information by preprocessing to accelerate the distance queries.
From the perspective of accuracy, distance oracles can be classified into approximate distance oracles (ADO) and exact distance oracles (EDO). ADOs may not answer the accurate distance but are cheaper in preprocessing time, query time, and index space, and thus scale to large graphs. EDOs always give the exact distance but can be more expensive to compute. Therefore, EDOs are usually used in applications that are on small graphs but more sensitive to accuracy. In this section, we will introduce how to apply our cluster-BFS to the two existing distance oracles: a 2-hop labeling-based EDO and a landmark labeling-based ADO. Since both applications work on undirected graphs, we assume the graph to be undirected in this section. Our cluster-BFS algorithm works for general directed graphs.
4.1   An EDO based on 2-Hop Labeling
Here, we consider the EDO constructed by Akiba et al. [4], called 2-hop labeling, and we apply the idea of C-BFS to this algorithm. Their original algorithm is sequential, and we refer to it as the AIY algorithm. We can replace the component for 2-hop labeling in their algorithm with our parallel C-BFS to achieve better performance. For completeness, we describe their algorithm in section D.1. In the experiments, we compared our parallel C-BFS with the component of C-BFS in their sequential code. We also parallelized the entire algorithm for 2-hop labeling using our parallel C-BFS, and present a comparison in section D.2.
1 The algorithm maintains $L[\cdot][\cdot]$ as the index with size $n\times|S|$. $L[v][i]$ is the distance between vertex $v$ and landmark $h_{i}$, initialized as $\infty$
2 Function Construct_Index$(G,H)$ // $G=(V,E)$
       // Each $h_{i}\in H$ is a vertex in the plain LL, and is a cluster in C-BFS-based LL
3       for $h_{i}\in H$  do
             // In our algorithm, we replace BFS with C-BFS
4             $t[1..n]\leftarrow\textsc{BFS}(G,h_{i})$
5             foreach $v\in V$ do  $L[v][i]\leftarrow t[v]$
6
7      return $L$
8Function Query$(u,v)$
9       $ans\leftarrow\infty$ // the answer of the query
10       for $i\leftarrow 0$ to $|H|-1$ do
             // Our algorithm computes the shortest distance via all vertices in all clusters
[…]
13      return ans
Algorithm 2 Framework of Landmark Labeling
4.2   An ADO based on Landmark Labeling
[…]
Landmark labeling (LL) is one of the widely-used approaches of ADOs and probably the simplest. The basic idea is to select a subset $H$ of vertices as landmarks, and precompute the distances between each landmark $h\in H$ and all the vertices $u\in V$. When the distance between two vertices, $u$ and $v$, is queried, $\mbox{query}(u,v)$ answers the minimum $\delta(u,h)+\delta(h,v)$ over all the landmarks $h\in H$ as an estimation. We show the high-level idea about LL in Alg. 2.
Generally, the distortion of a query depends on how far the actual shortest paths are from their nearest landmark. If the actual shortest paths contain any landmark vertex, the query answer is equal to the true distance. Therefore, adding more landmarks can decrease the distortion, but it also needs more space and time to store and compute the index. In this paper, we propose to use C-BFS to optimize LL. In particular, we select clusters of vertices as landmarks instead of single vertices. As discussed, a C-BFS on a cluster of size $O(\log n)$ has costs (both time and space) asymptotically the same as running BFS on one vertex. In this way, we can select $w$ times more landmarks with asymptotically the same cost as the plain LL. We note that the quality of the $w$ landmarks in one cluster may not be as good as choosing them independently, since they are highly correlated. However, we experimentally observe that with the same memory limit, using cluster-BFS in landmark LL significantly improves the performance both in distortion and preprocessing time (see section 5.3).
To apply C-BFS to LL, we need two subroutines: 1) selecting landmarks in clusters with low distortion, and 2) answering the queries from the cluster distance vectors computed by C-BFS.
Selecting Landmarks in Clusters.   The landmark selection is crucial in LL as it affects the query quality. A typical way is to prioritize vertices with high degrees [33, 38, 3]. We also employ this approach. In this step, we aim to identify clusters with a specified size $k=w$ and diameter $d$. Within these constraints, the selection of landmarks should prioritize vertices with higher degrees. To find a cluster, we first identify the vertex with the highest degree. Among all its $\lfloor d/2\rfloor$-hop neighbors, we then select $k-1$ additional vertices with the highest degrees. For instance, if $k=64$ and $d=2$, we select the vertex with the highest degree and $63$ of its neighbors with the next highest degrees to create the first cluster. Once we select a cluster, we will mark all the vertices in the cluster, and will not select them again. We repeat this process until we select $r$ clusters (giving $rw$ landmarks in total). One can also select landmarks using other heuristics [38, 34]. After selecting clusters, we apply C-BFS to all selected clusters using Alg. 1. By doing this, we obtain the cluster distance vectors between each vertex and each cluster.
[…]
The problem boils down to finding $\delta(u,s)+\delta(s,v)$ for all sources $s$ in a given cluster $S$. Recall that for each cluster $S$, both vertices $u$ and $v$ have obtained their cluster distance vector from C-BFS, denoted as $\langle S_{u}[0..d],\Delta_{u}\rangle$ and $\langle S_{v}[0..d],\Delta_{v}\rangle$. The possible shortest distances passing through $S$ is in the range $[\Delta_{u}+\Delta_{v},\Delta_{u}+\Delta_{v}+2d]$. For two bit-subset, ${S}_{u}[i]$ and ${S}_{v}[j]$, if their intersection is not empty, it means there is a path connecting $u$ and $v$ with distance $\Delta_{u}+\Delta_{v}+i+j$ passing through any source vertex in the intersection. We check the intersections from the lowest possible distances (e.g., ${S}_{u}[0]\cap{S}_{v}[0]$) to higher possible distances, until we find a nonempty intersection, and return their distance sum. The complexity of the query grows in a quadratic manner as $d$ grows. For the simplest case, $d=2$, we only need to check three intersections for each cluster to get the answer.
We have shown another optimization, bidirectional searching, for answering queries that can reduce distortion without much more overhead in querying time. Since this optimization is independent with C-BFS itself, due to the page limit, we put both the description and experiments in Sec. B.
5   Experiments
Setup.   We run our experiments on a 96-core (192 hyperthreads) machine with four Intel Xeon Gold 6252 CPUs and 1.5 TB of main memory. We implemented all algorithms in C++ using ParlayLib [9] for fork-join parallelism and parallel primitives (e.g., sorting). We use numactl -i all for parallel tests to interleave the memory pages across CPUs in a round-robin fashion.
We tested 18 undirected graphs, which are either social or web graphs with low diameters. Graph information is given in Tab. 2. All graphs are from commonly used open-source graph datasets [31, 18, 17, 43]. When comparing the average running times or speedups across all the graphs, we use the geometric mean.
Baseline Algorithms.   We compare our algorithm to two existing implementations: 1) Ligra, the parallel BFS in the Ligra [45] library (only using thread-level parallelism), and 2) AIY, which is the C-BFS component from Akiba et al. [4] (sequential, only using bit-level parallelism). We also compared AIY for the 2-hop distance oracle as one of the applications.
5.1   Microbenchmarks for Cluster-BFS
Graph Information Seq-BFS Related Work Parallel C-BFS Self-Speedup
Dataset $n$ $m$ Notes Time(s) AIY Ligra Final Time(s) C-BFS Ligra
EP 75.9K 811K Epinions1 [5] 0.18 20.6$\times$ 4.02$\times$ 102$\times$ 0.002 4.39$\times$ 9.44$\times$
SLDT 77.4K 938K Slashdot [32] 0.21 18.8$\times$ 3.91$\times$ 94.1$\times$ 0.002 3.47$\times$ 9.55$\times$
DBLP 317K 2.10M DBLP [55] 0.77 20.3$\times$ 6.22$\times$ 183$\times$ 0.004 10.2$\times$ 17.1$\times$
YT 1.13M 5.98M com-youtube [55] 3.30 22.9$\times$ 17.8$\times$ 445$\times$ 0.007 20.6$\times$ 31.6$\times$
SK 1.69M 22.2M skitter [43, 19] 6.56 21.0$\times$ 30.4$\times$ 496$\times$ 0.013 26.7$\times$ 33.1$\times$
IN04 1.38M 27.6M in_2004 [18, 17] 4.08 20.9$\times$ 4.00$\times$ 171$\times$ 0.024 10.1$\times$ 17.8$\times$
LJ 4.85M 85.7M soc-LiveJournal1 [5] 39.8 23.7$\times$ 61.8$\times$ 1017$\times$ 0.039 47.7$\times$ 53.9$\times$
HW 1.07M 112M hollywood_2009 [43, 18] 18.7 20.9$\times$ 89.7$\times$ 928$\times$ 0.020 32.4$\times$ 48.7$\times$
FBUU 58.8M 184M socfb-uci-uni [43, 50, 41] 268 32.0$\times$ 49.6$\times$ 973$\times$ 0.276 54.4$\times$ 52.8$\times$
FBKN 59.2M 185M socfb-konect [43, 50, 41] 176 27.9$\times$ 38.8$\times$ 712$\times$ 0.247 53.1$\times$ 51.8$\times$
OK 3.07M 234M com-orkut [55] 61.6 19.8$\times$ 102$\times$ 1119$\times$ 0.055 49.0$\times$ 65.4$\times$
INDO 7.41M 301M indochina [43, 17, 15] 38.8 21.9$\times$ 12.4$\times$ 452$\times$ 0.086 25.7$\times$ 35.9$\times$
EU 11.3M 521M eu-2015-host [18, 17, 16] 119 23.9$\times$ 26.6$\times$ 821$\times$ 0.145 18.5$\times$ 41.3$\times$
UK 18.5M 523M uk-2002 [18, 17] 91.8 22.7$\times$ 30.7$\times$ 687$\times$ 0.134 42.1$\times$ 46.7$\times$
AR 22.7M 1.11B arabic [18, 17] 147 22.5$\times$ 10.7$\times$ 461$\times$ 0.319 18.0$\times$ 33.8$\times$
TW 41.7M 2.41B Twitter [30] 861 20.6$\times$ 157$\times$ 856$\times$ 1.006 56.3$\times$ 60.2$\times$
FT 65.6M 3.61B Friendster [55] 2084 20.4$\times$ 187$\times$ 813$\times$ 2.563 59.4$\times$ 64.6$\times$
SD 89.2M 3.88B sd_arc [36] 1898 25.0$\times$ 80.3$\times$ 945$\times$ 2.008 55.7$\times$ 62.5$\times$
GeoMean 32.0 22.4$\times$ 27.0$\times$ 500$\times$ 0.064 24.8$\times$ 35.4$\times$
Table 2: Tested graphs and microbenchmarks on different BFS algorithms from a cluster of vertices with size 64. The numbers endup with ‘$\times$’ are speedups, higher is better. Others are running time, lower is better. The columns “AIY”, “Ligra” in related work and “Final” show the speedup over the “Seq-BFS”. “AIY” is referred to sequential C-BFS from [3], “Ligra” is referred to parallel single BFS [45], and “Final” is referred to our parallel C-BFS. The “self-speedup” is the speedup running the algorithm in parallel over running it in sequential.
Figure 3: Speedup of parallel Ligra BFSs and parallel C-BFS over the standard sequential BFS on cluster with size 64. $y$-axis is the speedup over sequential regular BFS in log-scale, higher is better. Each group of bars represents a graph, except the last group, which represents the average across all graphs. The numbers on the bar are the speedup of parallel algorithms over the standard sequential algorithm.
We start with testing our cluster-BFS (C-BFS) as a building block. We first test the simplest case where $d=2$ and $k=w=64$, where each cluster is a star (a vertex and its 63 neighbors). Our experiment runs on 10 different clusters on each graph and the average is reported. We present a detailed experimental study for this simple case because one of our baselines, AIY, only supports sources as star-shaped clusters. Another reason is that in previous work (as well as new results in this paper), we observed that using $d=2$ gives the best overall performance for the two applications discussed in this paper. We present some studies about varying $d$ at the end of this subsection.
Recall that our new C-BFS benefits from two aspects: 1) using bit-level parallelism with the idea of clustering to compute the results from $O(w)$ sources simultaneously, and 2) using thread-level parallelism and known parallel techniques for optimizing BFS (e.g., directional optimization). In our test, we choose ten different clusters and report the average running time of them, as well as the plain sequential BFS as the simplest baseline. The plain sequential BFS processes all 64 sources independently, and the running time is in the column “Seq-BFS time” in Tab. 2. Our final running time of C-BFS using all techniques is provided in the column “Par-Time (s)” in Tab. 2. To evaluate the performance gain by both techniques, we compared C-BFS with both Ligra and AIY. AIY only supports C-BFS on star-shaped clusters. The column “AIY” and “Ligra” in Tab. 2 provide the speedups over the plain sequential BFS. To better illustrate the results, we show the speedups relative to the plain version “Seq-BFS” in Fig. 3. Essentially, the column “AIY” means the speedup that can be achieved by applying bit-level parallelism on a cluster-BFS in the sequential setting. Similarly, the column “Ligra” provides the speedup that can be achieved by applying thread-level parallelism for running $k=64$ regular (non-cluster) BFS.
As shown in the column “AIY”, using clusters and bit-parallelism brings up 18.8–32.0$\times$ improvement, which is uniform on different graphs. Note that here, we have $k=64$ sources, so the maximum speedup can be 64$\times$. Of course, since C-BFS is more complicated than the plain BFS, there are some constant overheads, leading to an average of 22.4$\times$ speedup in a sequential setting.
For applying thread-level parallelism, the improvement on different graphs varies greatly, from 3.91$\times$ (on smallest graphs) to up to 187$\times$ on a 96-core machine with hyperthreads. The benefit on certain graphs (e.g., OK, TW, and FT) is significant, which is over 100$\times$. One reason for the difference in improvement is the directional optimization. On some dense graphs, directional optimization may exist early in the backward traverse and thus save work (see Sec. A for more details). This is consistent with the observations in existing work [45, 8]. On average, effectively utilizing parallelism gives 27.0$\times$ speedup over plain sequential BFS.
The columns “Par-Time(s)” and “Final” show our parallel C-BFS running time and overall improvement over “Seq-BFS”. Our algorithm combines the strengths of both bit-level and thread-level parallelism. Our solution is always better than any of the baselines. Compared to the plain sequential baseline, the improvement is 94.1–1119$\times$, and 500$\times$ on average.
For comparing the techniques and improvements of all baselines, we show a summary figure in Fig. 1. The time and speedup numbers are average on all tested graphs. Compared to Ligra, our algorithm improves the performance by 18.5$\times$ by utilizing clusters and bit-level parallelism. Compared to AIY, our algorithm improves the performance by 22.3$\times$ by utilizing thread-level parallelism. This indicates that our combination of bit- and thread-level parallelism works very well in synergy. Each of them still (almost) fully contributes to the performance, achieving the same level of improvement as when used independently. Therefore, we believe our work on an efficient implementation combining thread- and bit-level parallelism fills the gap in the existing study of both C-BFS and parallel BFS.
Self-relative Speedup.   In addition to the aforementioned set of baselines, we further tested C-BFS in the sequential setting to study the self-relative speedup (in column “Self-Spd.”). The speedup numbers are from 9.44$\times$ (on smallest graphs) to more than 40$\times$ (on most large graphs). In summary, the self-speedup of applying thread-level parallelism is 35.4$\times$ on average.
Figure 4: The scalability curve on different number of processors for C-BFS. The y-axis is the self speedup. The C-BFS running on one core is always 1. The x-axis is the number of cores. 96h represents 96 cores with hyperthreads.
Scalability on Different Number of Cores.   In addition to the overall self-speedup on 96 cores with hyperthreads, the scalability curve, as shown in Fig. 4, presents the self-speedup results across different core configurations. The curve demostrates that our algorithm achieves nearly linear speedup for most of the graphs, indicating efficient parallel scalability. One exception is the DBLP graph, which deviates from this pattern due to its smaller size, resulting in a running time of less than 0.02 seconds.
Influence of Cluster Diameter $d$ in C-BFS.   There are no existing implementations supporting C-BFS with general $d$. Recall that our C-BFS supports general clusters with diameter $d$ instead of star-shaped clusters (a vertex and its neighbors, where $d=2$). As shown in Thm. 3.1, the work is proportional to $d$. We tested different $d$ from $2$ to $6$ on all the graphs. We choose 10 representative graphs and show the C-BFS running time on clusters with different $d$ in Fig. 5. The full running time is shown in Tab. 4 in section 5.4. The running time increases as $d$ grows, especially when $d$ is small. A large $d$ allows for better flexibility for the shape of the cluster (the vertices can be further from each other), but significantly affects the performance. In section 5.4, we will further show that using a large $d$ in LL incurs overhead in space and time, and therefore, using $d=2$ is almost always more effective in that application.
Figure 5: The running time of C-BFS on various cluster diameter $d$. The $y$-axis shows the relative running time over $d=2$. The $x$-axis shows the cluster diameter $d$.
5.2   2-Hop Distance Oracle
As mentioned in Sec. 4, we can replace the C-BFS in Akiba et al. to accelerate their algorithm for an EDO. Due to the page limit, we present the results in section D.2. To do this, we also need to parallelize the other parts in their algorithm. In a nutshell, our algorithm with thread-level parallelism accelerates their algorithm by 9–36$\times$, and can also process much larger graphs than they can do.
5.3   Approximate Landmark Labeling
Index Time (s) $\epsilon$ (%)
Data Plain $w=64$ $w=8$ Plain $w=64$ $w=8$
EP 1.26 0.02 0.08 0.4 0.1 0.1
SLDT 1.15 0.02 0.07 0.7 0.1 0.1
DBLP 3.57 0.08 0.25 2.5 2.2 1.0
YT 9.22 0.23 0.59 0.3 0.3 0.1
SK 13.4 0.55 1.77 1.4 0.7 0.4
IN04 20.0 0.96 3.88 2.1 1.9 0.9
LJ 36.2 1.72 5.63 5.0 4.3 3.5
HW 12.4 0.93 4.10 10.6 5.6 7.1
FBUU 138 11.3 27.0 6.2 11.9 6.9
FBKN 127 10.5 24.9 6.2 11.9 6.9
OK 26.3 2.87 10.1 8.7 7.7 7.3
INDO 83.2 5.44 29.8 3.1 1.5 1.3
EU 87.3 7.01 34.9 2.6 1.3 1.7
UK 80.4 8.28 38.8 3.9 4.9 3.1
AR 148 17.6 86.8 2.6 4.0 2.2
TW 112 31.0 99.3 1.5 1.4 1.1
FT 251 61.1 193 16.8 12.4 12.8
SD 318 75.6 255 0.6 0.3 0.3
Table 3: The index construction time, (1$+\epsilon$) distortion, and query time for ADO based on landmark labeling. The “Plain” is the plain LL algorithm that each landmark is a single vertex. The “$w=64$” and “$w=8$” C-BFS-based LL that landmarks are in clusters with size $w$. The memory budget is 1024 bytes per vertex. For both index time and $\epsilon$, lower is better.
Figure 6: Tradeoffs between index size and distortion/construction time The $x$-axis is the memory limits per vertex in bytes, and is in log-scale. The $y$-axis on the left shows the ($1+\epsilon$) distortion. The $y$-axis on the right shows the preprocessing time. For both preprocessing time and distortion, lower is better. For the algorithms compared here, ‘plain’ is the regular LL, others are the C-BFS-based LL that choose clusters with size $w$ as landmarks.
We now show how C-BFS can significantly improve the landmark labeling (LL) approach regarding both running time and accuracy. In general, more landmarks will lead to better accuracy for the distance queries, as the landmarks are more likely to be on or close to the shortest path between the queried vertices. Hence, by using the clusters as landmarks, we can drastically increase the number of landmarks by a factor of $w$ with $O(d)$ overhead in time and space. For simplicity, we start with the clusters with diameter $d=2$, and we will discuss the clusters for $d>2$ later. Following the optimization mentioned in section 4.2, we use each cluster as a vertex and its $w-1$ neighbors, and only store a distance and $d$ bit-subsets in the cluster distance vector.
We study the effectiveness of our approach by comparing our C-BFS-based LL with a standard solution where each landmark is one vertex. We limit the total memory usage for both algorithms to be $t$ bytes per vertex for each graph, and construct an LL-based index within this memory budget. Note that the memory usage for C-BFS requires storing the following information per cluster: the distance (1 byte), and $d$ bit-subsets ($w/8$ bytes each), while the memory usage for a regular LL is one byte (the distance) per vertex. For example, C-BFS with $w=64$ and $d=2$ needs $17$ bytes to store a cluster distance vector, and C-BFS with $w=8$ and $d=2$ only needs $3$ bytes. Thus, the memory usage depends on the word size and number of clusters we choose. For fair comparisons, we fix the memory usage per vertex in the index for different baselines. With the same memory budget, a larger $w$ results in fewer (independent) clusters, thus usually allowing for faster preprocessing time. As discussed in section 4.2, the landmarks in the same cluster are highly correlated to each other, and may not bring the same benefit as independent ones. Namely, a larger $w$ may also result in less accuracy.
In this experiment, we tested on different $w$ from $\{8,16,32,64\}$ and $d=2$. In Tab. 3, we show the two extremes of using $w=64$ and $w=8$. The full information of all tested $w$ can be found in Sec. C.
We show the full result of using memory limit as $t=1024$ bytes per vertex in Tab. 3, where “plain” refers to using simple parallel BFS, “$w=64$” is to use C-BFS with $w=64$ (17 bytes per cluster), and “$w=8$” is C-BFS with $w=8$ (3 bytes per cluster). With the memory limit of 1024 bytes per vertex, we can choose 1024 landmarks for regular LL, 341 clusters for CC8, or 60 clusters for CC64. For regular LL, each landmarks are chosen based on prioritizing the high-degree vertices. We compare the index time and the error $\epsilon$ shown in percentage. It means that the ADO has $(1+\epsilon)$ distortion (defined in section 4.2). We take the average of 100,000 pairs to estimate the error on all the graphs, except for the largest two graphs FT and SD. We only compute the distance of 10,000 pairs on these two graphs, since generating the ground truth is expensive on large graphs.
On all 18 tested graphs, C-BFS with $w=64$ always gives a lower running time. When $w=8$, the running time is 2.4–5.5$\times$ higher than $w=64$, but still mostly faster than the plain version. With the same memory budget, $w=64$ roughly processes 5.68$\times$ fewer clusters than $w=8$. Similarly, comparing $w=64$ and the plain version, the number of (clustered) BFSs performed by $w=64$ is 17×17$\times$17 × fewer than the number of (single) BFSs by regular LL. The running time can be up to 53×53$\times$53 × faster, but on average, it is around 10$\times$—each C-BFS is still more expensive than a single BFS, but the numbers indicate that the overhead is small.
Regarding distortions, $w=8$ generally gives better accuracy than $w=64$, but it can also be worse in several instances. That is because $w=64$ selects more landmarks than $w=8$ (3840 vs. 2728) but fewer independent clusters (60 vs. 341). Therefore, the loss of using fewer clusters may or may not be compensated by more (correlated) landmarks. Empirically, the results still suggest that $w=8$ gives overall better accuracy, as it is more accurate on 11 out of 18 tested graphs. Both $w=8$ and $w=64$ are more accurate than the plain version on at least 16 out of 18 graphs. This indicates that the increased number of landmarks, although less independent, still positively affects the accuracy on most of the 18 scale-free networks we tested here.
It is worth noting that the query times for all graphs are similar, and we show the average (geometric mean) query time at the last line in Tab. 3. This is because the query time is completely determined by the number of landmarks, instead of the graph size.
The experimental results suggest that, if the primary objective is to reduce preprocessing time, using C-BFS with $w=64$ will always give a more efficient version than the plain version. The precision is also superior in most cases. When achieving better precision is the top priority, C-BFS with a smaller $w$ can significantly reduce the distortion. When using $w=8$, the distortion is almost always better than both $w=64$ and the plain version, and the running time also improves the plain version on most of the graphs.
We also show the trade-off between preprocessing time and distortion in Fig. 6 on three representative graphs, particularly including those that $w=8$ and $64$ may perform poorly on. We include results using $w=16$ and $w=32$, with different memory budgets ($t$ bytes per vertex, shown in the $x$-axis). In general, it is still true that large $w$ results in faster preprocessing but larger distortion. The red line shows the baseline of regular LL. For both preprocessing time and distortion, lower is better. Overall, using $w=16$ or $w=32$ provides a compromise for both time and distortion, and can be a more stable choice across all graphs.
For both preprocessing time and distortion, using C-BFS provides a significant improvement on almost all graphs. This verifies the effectiveness of our C-BFS to achieve a practical distance oracle.
C-BFS with Diameter $d>2$.   Unlike the (sequential) AIY algorithm that can only apply to $d=2$ case (a vertex and its neighbors), our C-BFS is general and works for any given $d$. Hence, it is interesting to understand how the performance and quality are affected by varying $d$. Note that larger $d$ gives us flexibility in selecting the sources. It can generally reduce the correlations between the sources, and even sometimes enable a larger $k$ if no vertices in the graph have large degrees (although this is rare in the scale-free networks tested in this paper). Therefore, we tested the time and distortion of LL on clusters with larger $d$ and present the results in section 5.4. The takeaway is that, although larger $d$ may provide higher source quality, the space and time consumption are also linearly proportional to $d$. Hence, given the same memory limit, choosing clusters with $d>2$ does not help with decreasing distortion (other than the graph OK, where $d=3$ improves the distortion over $d=2$ by $3\%$). However, generally, the faster running time is due to fewer clusters that can be selected for the same memory budget.
5.4   Further Study on the $d>2$ Case in Approximate Distance Oracle
Here, we provide additional information for the $d>2$ case for the landmark labeling, and continue the discussion from section 5.3. Recall $d$ is the diameter of clusters. We first study its influence in C-BFS. Then, we study the influence of $d$ in the application Landmark Labeling introduced in in section 4.2.
The Performance of C-BFS on Different $d$.   According to Thm. 3.1, the work of C-BFS is proportional to $d$. In order to know how to choose proper $d$ in different applications, we need to first know how different $d$ affect the running time of C-BFS in practice. We tested the running time of C-BFS with clusters of different diameter $d$. The results are shown in Tab. 4. The running time increases as $d$ grows larger, as shown in Tab. 4, and space usage ($O(d)$ space to store the distances of a cluster to a vertex) also increases. When applying C-BFS to other applications, we need to weigh the benefits brought by more general clusters and the overhead on time and space costs. Tab. 4 provides a reference for overhead on running time.
Data $d=2$ $d=3$ $d=4$ $d=5$ $d=6$
EP 0.002 0.002 0.003 0.003 0.003
SLDT 0.002 0.002 0.002 0.002 0.003
DBLP 0.004 0.005 0.005 0.007 0.008
YT 0.007 0.011 0.012 0.014 0.015
SK 0.013 0.017 0.020 0.023 0.026
IN04 0.024 0.027 0.029 0.032 0.033
LJ 0.039 0.053 0.065 0.076 0.082
HW 0.020 0.028 0.037 0.042 0.045
FBUU 0.276 0.368 0.451 0.585 0.654
FBKN 0.247 0.318 0.388 0.501 0.564
OK 0.055 0.082 0.105 0.114 0.121
INDO 0.086 0.114 0.134 0.159 0.176
EU 0.145 0.195 0.239 0.284 0.290
UK 0.134 0.186 0.229 0.265 0.292
AR 0.319 0.420 0.485 0.581 0.647
TW 1.006 1.723 2.284 2.494 2.539
FT 2.563 4.412 5.880 6.594 6.815
SD 2.008 3.305 4.280 4.970 5.128
GeoMean 0.064 0.086 0.104 0.120 0.131
Table 4: The parallel C-BFS time (seconds) for one cluster with size 64 on different cluster diameter $d$.
The Performance of Different $d$ on Landmark Labeling.   Our C-BFS is the first implementation of C-BFS that supports general clusters. It gives us a chance to study the performance and quality of LL for clusters with larger $d$. Clusters with larger $d$ have fewer correlations for vertices in the same cluster, but the computational cost for C-BFS also increases. We are interested in whether it is worth using clusters with larger $d$ in the LL application.
We show the full result of using memory limit as $t=1024$ bytes per vertex and word size $w=64$ in Tab. 5, where “plain” refers to using simple parallel BFS (1 bytes per landmark), “$d=2$” is to use C-BFS with $d=2$ (17 bytes per cluster), “$d=3$” is C-BFS with $d=3$ (25 bytes per cluster) and “$d=4$” is C-BFS with $d=4$ (33 bytes per cluster). With memory limits 1024 bytes per vertex, we can choose 1024 landmarks for regular LL, 60 clusters for $d=2$, 40 clusters for $d=3$ , and 31 clusters for $d=4$. Landmarks are chosen based on prioritizing the high-degree vertices, similar to the setting previously.
The construction time is reversely proportional to $d$, since larger $d$ leads to fewer sources and clusters. However, for distortion, other than the graph OK, larger $d$ always leads to lower accuracy. For the graph OK, the improvement is only about $3\%$. In conclusion, since for distance oracles, generally the space usage and accuracy are the most crucial—given the same memory budget, choosing clusters from diameter $d>2$ does not help with accuracy on most graphs. However, since our C-BFS algorithm is highly parallel and efficient, it provides opportunities for researchers in the future to study other applications on whether more general sources with $d>2$ can be more effective than star-size ones with $d=2$.
Construction Time (s) Distortion $\epsilon$ (%)
Data Plain $d=2$ $d=3$ $d=4$ Plain $d=2$ $d=3$ $d=4$
EP 1.26 0.03 0.02 0.02 0.4 0.1 0.2 0.2
SLDT 1.15 0.02 0.02 0.02 0.7 0.1 0.4 0.6
DBLP 3.57 0.09 0.07 0.06 2.5 2.2 3.4 4.0
YT 9.22 0.23 0.19 0.17 0.3 0.3 0.6 0.8
SK 13.4 0.58 0.41 0.37 1.4 0.7 1.5 2.1
IN04 20.0 1.06 0.70 0.58 2.1 1.9 2.4 2.6
LJ 36.2 1.79 1.41 1.26 5.0 4.3 5.5 6.0
HW 12.4 0.99 0.75 0.63 10.6 5.6 6.5 7.0
FBUU 138 11.7 9.31 8.26 6.2 11.9 13.4 15.9
FBKN 127 10.8 8.82 7.52 6.2 11.9 13.5 16.0
OK 26.3 3.05 2.65 2.32 8.7 7.7 7.5 7.6
INDO 83.2 6.09 3.92 3.40 3.1 1.5 2.2 2.6
EU 87.3 8.00 6.33 5.55 2.6 1.3 1.9 2.2
UK 80.4 9.06 6.63 5.90 3.9 4.9 5.3 5.6
AR 148 19.4 13.4 11.7 2.6 4.0 6.8 7.1
TW 112 31.0 27.8 23.9 1.5 1.4 1.5 1.8
FT 251 61.1 52.0 45.8 16.8 12.4 13.7 14.6
SD 318 75.6 61.2 53.1 0.6 0.3 0.8 0.9
Table 5: The Approximate Landmark Labeling Time and Distortion for cluster with size 64 and memory limits 1024 bytes per vertex on different cluster diameter $d$.
6   Related Work
This paper mainly focuses on scale-free (small-diameter) networks, and we refer the audience to an excellent survey [6] of algorithms for large-diameter graphs (e.g., road networks). In this paper, we discuss the approaches based on landmark labeling, and here, we review other approaches for distance queries. We first review the approximate solutions. The concept of approximate distance oracles (ADOs) was proposed by Thorup and Zwick [49], and has later been theoretically studied in dozens of papers. Practically, papers [38, 29, 24] discuss how to select the best “landmarks” for these type of sketch-based solutions, and showed that degrees or betweenesses are reasonable metrics. Other solutions include embedding-based solutions [57, 40, 44, 35] (embedding graph metric into other simpler ones such as Euclidean space), tree-based approaches [13, 56], and some recent attempts using deep learning [42]. Most of these approaches are more complicated than the landmark-based labeling mentioned in this paper.
Regarding the exact distance queries, Pruned Landmark Labeling (PLL) [3] is among the latest solutions for scale-free networks. Follow-up work [33] provided detailed improvements, but the lack of publicly available code prevents a direct comparison with our approach. Additionally, other techniques, such as contraction hierarchies (CH) [26] and transit nodes routing [7], focus more on large-diameter graphs like road networks.
7   Conclusion
In this paper, we present parallel implementations for cluster-BFS, which runs BFS from a cluster of vertices with diameter $d$. Our algorithm is work-efficient in theory, and also leads to high parallelism on low-diameter graphs as we tested in the experiment. We employ both bit-level and thread-level parallelism to optimize the performance. Both of them lead to significant speedup. Especially, we observed that bit-level and thread-level parallelism work well in synergy. We also show that the combination of the techniques also leads to performance improvement in two applications in distance oracles in multiple measurements of preprocessing time and accuracy, and allows our implementation to scale to much larger graphs than a sequential algorithm. Besides, our C-BFS is the first implementation that supports general clusters with diameter $d$ instead of star-shaped clusters, which give us a chance to study the benifits and overhead of choosing clusters with larger $d$ in C-BFS and in its applications.
A   Parallel BFS
We briefly review direction-optimizing parallel BFS since it is the state-of-the-art parallel BFS [25], and many of the ideas are also used in our parallel cluster-BFS. We present the high-level idea of BFS from a single source $s\in V$ in Alg. 3. The algorithm maintains a frontier of vertices to explore in each round, starting from the source, and the algorithm finishes in $D$ rounds. It also maintains the distance from the source to each vertex in array $\delta$, initialized as infinity except for the source. In round $i$, the algorithm processes the current frontier $\mathcal{F}_{i}$, adding their out neighbors to the next frontier $\mathcal{F}_{i+1}$ if the neighbor has not been visited. If multiple vertices in $\mathcal{F}_{i}$ attempt to add the same vertex to $\mathcal{F}_{i+1}$ to the next frontier, a compare_and_swap is used to ensure that only one can succeed in updating the distance from $\infty$ to the current distance.
Following the existing graph processing library Ligra [45, 46], we use the EdgeMap framework with the direction-optimizing parallel BFS (see Alg. 3 and 4). It maps a subset of vertices (current frontier) to another subset of vertices (next frontier) by applying a given function to the out-edges from the current frontier. EdgeMap requires two user-defined functions, Cond_F and Edge_F. Cond_F$(v)$ is a function to indicate whether the vertex $v$ needs further processing. In BFS, it checks whether the vertex has not been visited—i.e., whether its distance is still $\infty$ (Alg. 3 in Alg. 3). Edge_F$(u,v)$ is a function for edge $(u,v)$, which processes the edge, and returns a boolean value indicating whether $v$ should be added to the next frontier by $u$. In BFS, it sets the distance of $v$ to one plus the distance of $u$. When multiple vertices want to add $v$ at the same time, $u$ should add $v$ iff. it wins compare_and_swap (Alg. 3 in Alg. 3).
The idea of direction optimization means to implement EdgeMap in two different “modes”: EdgeMap-Sparse (forward) and EdgeMap-Dense (backward), as shown in Alg. 4. In the sparse (forward) mode, we start with the current frontier and consider its out-neighbors as mentioned above. However, if the number of out-neighbors of the frontier is sufficiently large (i.e., close to $n$), it can be more efficient to use the dense (backward) mode instead, where all unvisited vertices look at all their in-neighbors to see if there is any on the frontier. With a large frontier, the dense mode can avoid costly atomic operations and make better use of cache locality, giving better performance. EdgeMap is a building block for parallel BFS and many other vertex-based graph algorithms. Our cluster-BFS also uses EdgeMap with the directional optimization.
Input: A graph $G=(V,E)$ and a source $s\in V$
1 $\delta\leftarrow\{\infty,...,\infty\}$
2 $\delta[s]\leftarrow 0$
3 $\mathcal{F}_{0}\leftarrow\{s\}$
4 $i\leftarrow 0$
[…]
7 while $\mathcal{F}_{i}\neq\emptyset$ do
[…]
9       $i\leftarrow i+1$
return $\delta$
Algorithm 3 Framework of Parallel BFS
Input: A subset of vertices $\mathcal{F}_{in}$, a condition function for vertex Cond_F, and a mapping function for edge Edge_F
Output: A subset of vertices $\mathcal{F}_{out}$
[…]
2       if $N^{+}\mathopen{}\mathclose{{}\left(\mathcal{F}_{in}}\right)$ is large then
3            return EdgeMap-Dense($\mathcal{F}_{in}$, Cond_F, Edge_F)
4      else
5            return EdgeMap-Sparse($\mathcal{F}_{in}$, Cond_F, Edge_F)
6Function EdgeMap-Sparse($\mathcal{F}_{in}$, Cond_F, Edge_F)
7       $\mathcal{F}_{out}=\emptyset$
8       ParallelForEach  $\{(u,v)~{}|~{}u\in\mathcal{F}_{in},v\in N^{+}(u)\}$ do
9             if Cond_F$(v)$ then
10                   if Edge_F$(u,v)$ then  $\mathcal{F}_{out}\leftarrow\mathcal{F}_{out}\cup\{v\}$
11
12      return $\mathcal{F}_{out}$
13Function EdgeMap-Dense($\mathcal{F}_{in}$, Cond_F, Edge_F)
14       $\mathcal{F}_{out}=\emptyset$
15       ParallelForEach $v\in V$ do
16             if Cond_F$(v)$ then
17                   for $u\in N^{-}(v)$ do
18                         if $u\in\mathcal{F}_{in}$ and Edge_F$(u,v)$ then
19                               $\mathcal{F}_{out}\leftarrow\mathcal{F}_{out}\cup\{v\}$
20                               break
21      return $\mathcal{F}_{out}$
Algorithm 4 Framework of EdgeMap
B   Applying Bi-Directional Search in the Queries
In the ADO based on landmark labeling, the distortion between two faraway vertices can be reasonably good, but the returned distance of two nearby vertices may not be as accurate [4]. This is because when the shortest path between $u$ and $v$ is short, it is less likely to pass a landmark, or to be close to a landmark. To further improve the accuracy, several techniques were proposed [29, 51, 39]. They typically store shortest-path trees rooted at the landmarks instead of just storing distances, and finding loops or shortcuts on the trees during the query to reduce the distortion. While they improve the accuracy, the query time becomes much slower. In our implementation, we add a fixed-size bi-directional search between queried vertices $u$ and $v$ before we run the actual query using landmarks. In particular, we will search $\tau$ vertices from each side by a BFS order, and take an intersection to find the minimum distance among them. Conceptually, the vertices encountered in the search can be viewed as landmarks generated on the fly. The intuition is that, for two nearby vertices, a bidirectional search should take a short time but give an exact distance with no distortion. In our experiments, when choosing a proper search size $\tau$, this optimization greatly improved query quality with a small overhead in query time.
Figure 7: Tradeoffs between local search size and distortion/query time. With a memory limit of 1024 bytes per vertex for the index, we vary the local search size from 128 to 2048 to see the influence on distortion and query time. The $x$-axis is the local search size, which is the total number of vertices explored during the bi-directional BFS from two queried vertices. The $y$-axis on the left is the ($1+\epsilon$) distortion, corresponding to the solid lines in the figures. The $y$-axis on the right is the parallel query time for $10^{4}$ queries in milliseconds, corresponding to the dashed lines in the figures. For both distortion and preprocessing time, lower is better. For algorithms compared here, ‘Plain’ is the regular LL, and others are C-BFS-based LL with cluster size $w$.
Performance Study of Query Optimizations.   As we mentioned, using a bidirectional search may help improve the accuracy for close-by vertices. We test different local search sizes and show the improvement in distortion and their impact on query time in Fig. 7. For almost all cases, using a local search size of 512 or more, we can see a clear improvement in distortion. However, with the local search size reaches 2048, the query time increases dramatically.
C   Approximate Landmark Labeling Full Information
In the paper, due to page limits, we only show the two extremes $w=64$ and $w=8$. We provide the entire experimental results for all the tested $w$, and their query time in Tab. 6.
Index Time(s) $\epsilon$(%) Query Time(ms)
Data Plain $w=64$ $w=32$ $w=16$ $w=8$ Plain $w=64$ $w=32$ $w=16$ $w=8$ Plain $w=64$ $w=32$ $w=16$ $w=8$
EP 1.26 0.02 0.04 0.05 0.08 0.4 0.1 0.1 0.1 0.1 2.4 2.1 2.0 1.7 2.4
SLDT 1.15 0.02 0.03 0.06 0.07 0.7 0.1 0.1 0.1 0.1 2.0 2.0 1.9 1.9 2.4
DBLP 3.57 0.08 0.11 0.16 0.25 2.5 2.2 1.4 1.1 1.0 2.0 1.8 1.9 1.9 2.5
YT 9.22 0.23 0.27 0.37 0.59 0.3 0.3 0.2 0.2 0.1 1.7 1.8 2.1 2.1 2.5
SK 13.4 0.55 0.81 1.14 1.77 1.4 0.7 0.5 0.4 0.4 1.7 1.8 2.0 2.0 2.7
IN04 20.0 0.96 1.84 2.38 3.88 2.1 1.9 1.3 1.1 0.9 1.7 1.8 2.2 2.0 2.1
LJ 36.2 1.72 2.48 3.49 5.63 5.0 4.3 3.7 3.6 3.5 1.8 1.8 2.0 2.0 2.3
HW 12.4 0.93 1.73 2.49 4.10 10.6 5.6 5.9 6.5 7.1 1.7 1.8 2.1 2.1 2.4
FBUU 138 11.3 13.1 17.8 27.0 6.2 11.9 9.9 8.1 6.9 1.8 1.8 2.0 1.9 2.2
FBKN 127 10.5 12.1 16.3 24.9 6.2 11.9 10.0 8.2 6.9 1.8 1.8 2.0 1.9 2.2
OK 26.3 2.87 4.53 6.43 10.1 8.7 7.7 7.6 7.3 7.3 1.8 1.8 2.0 2.0 2.3
INDO 83.2 5.44 11.1 18.0 29.8 3.1 1.5 1.3 1.2 1.3 1.8 1.8 2.0 1.9 2.0
EU 87.3 7.01 14.5 21.9 34.9 2.6 1.3 1.0 1.2 1.7 1.8 1.8 2.0 1.9 2.0
UK 80.4 8.28 15.8 22.1 38.8 3.9 4.9 3.8 4.1 3.1 1.8 1.8 2.0 1.9 2.1
AR 148 17.6 36.6 54.9 86.8 2.6 4.0 3.2 2.2 2.2 1.8 1.8 2.0 1.9 2.0
TW 112 31.0 47.0 62.7 99.3 1.5 1.4 1.2 1.1 1.1 1.8 1.8 1.9 1.9 2.3
FT 251 61.1 88.2 120 193 16.8 12.4 12.0 12.1 12.8 1.8 1.8 2.0 1.9 2.4
SD 318 75.6 125 163 255 0.6 0.3 0.3 0.3 0.3 1.8 1.8 1.9 1.9 2.4
Table 6: The index construction time, (1$+\epsilon$) distortion, and query time for ADO based on landmark labeling. The “Plain” is the normal LL algorithm in which each landmark is a single vertex. Others are C-BFS-based LL that landmarks are in clusters with size $w$. The memory budget is 1024 bytes per vertex. For both index time and $\epsilon$, lower is better.
D   Exact 2-Hop Distance Oracle
D.1   Algorithm Description of Exact 2-Hop Distance Oracle
Our second application is an exact distance oracle, which always answers the correct shortest distance in queries. Many EDOs are based on the idea of 2-hop cover [22, 21, 2], described as follows.
For each vertex $v$, 2-hop labeling methods select a subset of vertices $u\in V$ as hubs for $v$, and precompute their distance $\delta(u,v)$. We call the precomputed distances for a vertex $v$ as the label of $v$, and denote it as $L(v)$: a set of pairs $(u,\delta(u,v))$ for each of $v$’s hub $u$. Note that the hubs can be different for different vertices. Then, $query(s,t)$ finds the shortest distance passing through the intersection of their hubs, i.e., $\min\{\delta(s,v)+\delta(t,v)~{}|~{}(v,\cdot)\in L(s)\cap L(t)\}$. We call $L$ a 2-hop cover of $G$ if $query(s,t)$ can correctly answer the distance between any pair of vertices. Finding a small 2-hop cover efficiently is a long-standing challenge [22, 21, 2].
One of the state-of-the-art approaches is Pruned Landmark Labeling (PLL) [3]. Given a graph $G$ and a vertex order $v_{1},v_{2},\dots,v_{n}$, PLL runs BFSs (with pruning, introduced below) from vertices in order and construct the labels. PLL starts with an empty index $L_{0}$, where $L_{0}(v)=\emptyset$ for every $v\in V$. In round $i$, PLL conducts a BFS from vertex $v_{i}$, and adds distances from $v_{i}$ to labels of reached vertices, that is, $L_{i}(u)=L_{i-1}(u)\cup\{(v_{i},\delta(v_{i},u))\}$ for each $u$ that $v_{i}$ can reach. To minimize the size of the index, PLL prunes unnecessary labels and searches during each BFS from $v_{i}$: when $v_{i}$ visits $u$, PLL checks if the existing index can already report the distance between $v_{i}$ and $u$. If so, the BFS will skip $u$, and $v_{i}$ will not be added to the labels of $u$. PLL is proved to be a correct 2-hop cover, and the index constructed is minimal [3].
To further speed up both the preprocessing and querying, PLL [3] applies C-BFS. Instead of running pruned BFS, in the first several rounds, C-BFS is used (without any pruning), such that all sources in the clusters will be added to the labels of all vertices in $V$. The intuition is that, in the first several rounds, PLL can hardly prune any vertices since the index size is small. Therefore, we can ignore the pruning but use clusters to improve the performance. As a result, the entire PLL algorithm in [3] has two phases: 1) several rounds of C-BFSs and 2) pruned BFSs on the rest of the vertices. In our paper, we develop a parallel version for both C-BFS and pruned BFS to the sequential algorithm in [3].
Our Implementation.   We apply our parallel C-BFS to PLL, and also provide a parallel implementation for the pruned BFS. We directly apply our parallel C-BFS mentioned in section 3.2 with the optimizations mentioned in Sec. 4 to the first phase. Note that in the original PLL algorithm, the second phase incurs running pruned BFS from almost all vertices one by one. It is essential to parallelize this part to achieve high performance of the entire process of PLL. Ideally, we want to 1) run BFSs from multiple sources in parallel, but also 2) make them see as much of the index constructed by other vertices to enable effective pruning. To do this, we use a prefix-doubling-like scheme [11, 12, 47], which splits the vertices into batches of exponentially growing sizes, and we parallelize the BFSs within each batch. The batches will be executed one by one from the smallest one. We empirically set the size of the $i$-th batch as $200\times(1.5)^{i}$, and stop increasing the batch size when the batch size is large enough (1000 in our implementation) for sufficient parallelism. With both phases well-parallelized, our parallel version improves the preprocessing time of the original sequential code from [3] by up to 36.5$\times$, and it can process much larger graphs than the sequential algorithm.
D.2   Experiments on Exact 2-Hop Distance Oracle
avg. Index Running Time
$r$ Alg. labs Size C-BFS P-BFS Total
SK 64 AIY 123 2.68 32.9 265 299
Ours 126 2.71 1.05 15.0 16.2
Spd 31.3$\times$ 17.7$\times$ 18.5$\times$
HW 64 AIY 2237 12.2 103 11010 11115
Ours 2280 12.4 1.26 303 304
Spd 82.3$\times$ 36.4$\times$ 36.5$\times$
INDO 64 AIY 323 18.7 246 3740 4051
Ours 349 19.6 5.63 421 428
Spd 43.7$\times$ 9.02$\times$ 9.46$\times$
EU 64 Ours 944 61.0 9.92 1385 1396
LJ 512 Ours 2585 97.7 23.0 2718 2742
AR 256 Ours 989 197 72.7 7690 7767
OK 2048 Ours 6881 198 119 13407 13527
Table 7: Performance on an exact distance oracle based on pruned landmark labeling. “AIY”: the sequential implementation from  [3]. $r$: the number of clusters used in C-BFS. Index sizes are in GB. “C-BFS”: time for cluster-BFS. “P-BFS”: time for pruned BFS. “Spd”: speed-up of ours over AIY. For #labels/vertex, index size, and running time, lower is better. See more details in section D.1.
We also apply our C-BFS to an exact distance oracle using the pruned landmark labeling (PLL) described in section D.1. We follow the high-level idea in [3], which consists of a cluster-BFS (C-BFS) phase on $r$ clusters and a pruned BFS (P-BFS) phase on the rest of vertices in $V$. We compare our parallel implementation with the original sequential code provided in [3] (referred to as the AIY algorithm). Note that due to the need to report the exact distance, the index size is much larger than the ADOs reported in section 5.3. For the baseline algorithm (AIY), the largest graph it can process is INDO with 7.41M vertices and 301M edges. Because of better parallelism, our C-BFS can scale to much larger graphs. We show four graphs (EU, LJ, AR, OK) that can be processed by our parallel algorithm but not the sequential version. The largest graph includes 22.7M vertices and 1.11B edges.
To choose the parameter $r$ (number of cluster searches), for all graphs that have been tested in [3], we use the same value $r$ as they reported giving the almost best memory usage. For other graphs, we test a wide range of $r$ and present the overall best performance considering both space and preprocessing time. We present the parameter $r$ and running time for each graph in Tab. 7.
As mentioned, our algorithm may result in more labels (thus larger sizes) over the original AIY algorithm, because of running BFS in batches in the pruned BFS phase: the vertices in the same batch may not be able to see and use each others’ labels for pruning, and thus more labels may be added. In our results, such a loss is reasonably small. On all tested graphs, it is at most 8% of the number of labels and at most 5% more of the total index size.
For the running time, the time on P-BFS always dominates the cost, both in sequential and in parallel. Our algorithm parallelizes both steps well, with better speedup on the C-BFS phase. Note that although C-BFS is not the major cost of the sequential PLL, without parallelizing, it will make its cost comparable to or even larger than parallel P-BFS. Therefore, it is important to combine C-BFS with parallelism and make this part negligible in the parallel running time. In total, on the five small graphs, our algorithm achieves 6.3–36.5$\times$ speedup over the sequential AIY algorithm. The advantage of our algorithm is more significant with more expensive sequential running time.
On four larger graphs, our algorithm generates the index in 4 hours, and scales to graph AR with up to 22.7M vertices and 1.11B edges. Using a similar amount of time, the sequential AIY code can only process a much smaller graph (HW) with 1.07M vertices and 112M edges, about one order of magnitude smaller.